OpenStudy (arbyj):

how many micrograms are in 2.45 x 10^3 grams? the answer is 2.45 x 10^9 but how do i get it

OpenStudy (kl0723):

you have \[\frac{ 2.45x10^3 grams }{ }*\frac{ 1 microgram }{ 10^-6 grams }= 2.45x10^9 micrograms\]

OpenStudy (kl0723):

got it?

OpenStudy (arbyj):

is 10-6 10^6

OpenStudy (kl0723):

yes... sorry and then it goes to the top to become positive and multply with 10^3*10^6=10^9... so it makes 2.45x10^9

OpenStudy (arbyj):

how'd you get 10^6?

OpenStudy (kl0723):

you start with grams... so the trick is to multiply with grams on the bottom and the unit u want on top

OpenStudy (kl0723):

so you have 10^-6 grams on the bottom... the grams cancel with the other grams and you are left with micrograms... take 10^-6 to the top and it becomes positive: 10^6 ;)

OpenStudy (arbyj):

sorry, i meant to say how'd you get10^-6 lol

OpenStudy (arbyj):

@kl0723

OpenStudy (kl0723):

because the prefix micro has a value of 10^-6 base units

OpenStudy (arbyj):

@kl0723 that makes so much sense thank you